Fermi National Laboratory


Accelerator Update for December 9, 2005 - December 12, 2005

Friday December 9

The day shift began with the Tevatron (TeV) in access, with the Antiproton Source (Pbar) and Switchyard (SY) down due to power supply problems, and with MiniBooNE and NuMI taking beam.

Operations prepared NuMI for an access at 8:30 AM.

Power supply experts reported at 10:17 AM that they had completed their repairs on the TeV sector B2 VCB.

Power supply experts reported at 12:22 PM that they had completed their repairs on the Debuncher D:IB.

Operations had established beam to the Meson MTest and MCenter experiments by 1:21 PM.

A Main Injector (MI) expert began a study period at 2:04 PM.

Operations put the TeV through a dry squeeze at 2:05 PM.

Operations established beam to the TeV for studies at 2:43 PM.

NuMI resumed taking beam at 5:01 PM.

At 5:58 PM, Operations reported a problem with a Debuncher kicker (D:EKIK). Operators contacted power supply experts who, after 90 minutes, had accessed the tunnel and made a voltage adjustment, which resolved the problem.

Pbar experts began reverse protons studies at 8:09 PM.

TeV experts completed their studies at 8:45 PM. Operators put the separators into conditioning mode.

At 9:17 PM, Operations reported more problems with D:EKIK. Power supply experts accessed the tunnel, made further adjustments and reported the supply fixed at 10:24 PM.

Saturday December 10

The midnight shift began with TeV separators conditioning, with Pbar conducting studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

At 5:01 AM, a Booster RF station (BRF19) was found off with no alarm. Operations reset the station.

At 5:03 AM, there was a BRF8 airflow alarm.

At 5:11 AM, Operations reported that BRF8, BRF17, and BRF19 were all off and wouldn’t reset. Operators called in RF experts who had BRF8 repaired by the end of the shift.

Operations took the TeV B1 separators out of conditioning mode at 7:40 AM.

Operations turned off NuMI at 8:15 AM in preparation for an access.

At 9:10 AM, Operations turned off the Booster for twenty-six minutes while experts tied BRF17 into the system.

NuMI came out of access and resumed taking beam at 9:42 AM.

Operations turned off NuMI at 11:50 AM while experts repaired the Horn's chiller.

TeV studies ended at 6:27 PM. Operations put the separator back into conditioning mode. At 7:08 PM, a vertical separator sparked.

At 9:26 PM, the Meson MC7 enclosure interlocks dropped. Operators investigated and then called in interlock technicians to work on the problem.

Operations began putting the TeV through the first of a series of dry squeezes at 10:27 PM.

Sunday December 11

The midnight shift began with TeV separators conditioning, with the TeV going through dry squeezes, with Pbar conducting studies, with NuMI off due to a horn chiller problem, and with MiniBooNE, MTest, and MCenter taking beam.

At 2:29 AM, BRF1 and 2 tripped off for the first of two times in twenty minutes. Operators had to reset the stations manually after the second trip.

At 5:19 AM, the TeV C17 separator sparked.

Operations put the Collider Detector at Fermilab (CDF) and D-Zero (D0) into access at 6:01 AM.

CDF came out of access at 8:42 AM.

Pbar experts began tuning the Debuncher at 8:59 AM.

At 9:19 AM, a MI abort line vacuum bit came into alarm. MI experts told Operations that there should be no MI studies until this was fixed. Operations called a vacuum expert.

Pbar began stacking at 9:21 AM.

A vacuum expert suggested at 11:40 AM that Operations turn off the abort line vacuum pump to allow the pressure to equalize around the valve. Operations did this and it succeeded.

At 3:57 PM, Operations reported that a Debuncher RF cavity was sparking (DRF1-1), which was limiting the stacking rate. An expert came in to investigate, but then the problem disappeared.

At 5:19 PM, an Accumulator RF station (ARF1) tripped off for the first of four times in about an hour on a driver fault. It reset each time.

At 8:31 PM, Operations reported that ARF1 tripped off for the first of two times over the next three hours. It reset both times.

Operations prepared the Recycler to stash antiprotons at 8:51 PM. The stash ended and Pbar and the Meson experiments resumed taking beam around 10:21 PM.

Operations put the TeV into shot setup at 11:58 PM.

Monday December 12

The midnight shift began with TeV in shot setup, with NuMI off due to Horn chiller problems, and with MiniBooNE, MTest, and MCenter taking beam.

At 12:53 AM, the TeV suffered a sector B magnet quench (C:BQ9). No cryo systems quenched. This quadrupole that quenched was located in the B16 spool piece, which was not the one that was replaced.

Pbar and SY resumed taking beam at 1:20 AM.

At 2:19 AM, Operations reported that ARF1 tripped off for the first of three times in an hour. It reset each time.

Operations put the TeV through a dry squeeze at 4:01 AM.

Operations put the TeV into shot setup at 4:28 AM.

At 4:42 AM, a power supply (I:HV703) for a magnet located in the P1 transfer line tripped off on a ground fault indication. Operations called in an expert who checked the magnet and supply, but at 7:36 AM reported that he couldn’t find a problem.

The Plans for Monday

The plans are to increase the stacking rate, establish a store, and get back to regular stacking, stashing, and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/12/2005   email Fermilab